OpenStudy (anonymous):

Quadratic Formula to solve this x^4-12x^2-64=0

OpenStudy (anonymous):

@annas Do you know how to do this?

OpenStudy (cwrw238):

Let x^2 = y then y^2 - 12y - 64 = 0 you can solve this by factoring - no need to use the quadratic formula

OpenStudy (anonymous):

I have to use the quadratic formula if I want the credit. I know how to use the quadratic formula but since this isn't in the right form I can't.

OpenStudy (cwrw238):

(y - 16)(y + 4) = 0 y =16 or -4 thus x^2 = 16 or -4 if we only need real solution we can ignore -4 so we are left with x^2 = 16

OpenStudy (cwrw238):

oh ok useing quadratic formula y = -(-12) +/- sqrt((-12)^2 - 4*1*-64) / 2 = [12 +/- sqrt(144+256)] / 2 = [12 +/- sqrt400] / 2 = 6 + 10 or 6 - 10 = 16 or -4 now since y = x^2 x^2 = 16

OpenStudy (anonymous):

So x=-4, 4 or x=-2i or 2i

OpenStudy (cwrw238):

yes